英語 (English)

発信型英語 (Communicative English) (工 ((生 B)2 年))

ラックストンロバート · 非常勤講師 2 単位前期水 3 ・ 4

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】The purpose of this class is to improve English conversational

ability. 【授業の概要】 We will do textbook exercises, role-plays and skits. 【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 Students will increase their confidence about speaking English. 【授業の計画】

1. Introductions

2. Unit 1.1

3. Unit 1.2

4. Unit 2.1

5. Unit 3.1

6. Unit 3.2

7. Quiz 1 and Skit Preparation

8. Skit Day

9. Unit 4.1

10. Unit 4.2

11. Unit 5.1

12. Unit 5.2

13. Unit 6.1

14. Unit 6.2

15. Quiz 2 and Skit Preparation

16. Skit Day 【教科書】 Jack C. Richards: New Person to Person, Student Book 2 【参考書等】 [参考資料] 【成績評価の方法】 Quiz 1 25%; Skit 1 25%; Quiz 2 25%; Skit 1 25% 【再試験の有無】 [再評価] 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220911 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ ラックストン .

English

Communicative English (工 ((生 B)2 年))

Robert Luxton · Part-time Lecturer 2 units 前期水 3 ・ 4

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ The purpose of this class is to improve English conversational ability. Outline⟩ We will do textbook exercises, role-plays and skits. �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ Students will increase their confidence about speaking English.

Schedule⟩

1. Introductions

2. Unit 1.1

3. Unit 1.2

4. Unit 2.1

5. Unit 3.1

6. Unit 3.2

7. Quiz 1 and Skit Preparation

8. Skit Day

9. Unit 4.1

10. Unit 4.2

11. Unit 5.1

12. Unit 5.2

13. Unit 6.1

14. Unit 6.2

15. Quiz 2 and Skit Preparation

16. Skit Day Textbook⟩ Jack C. Richards: New Person to Person, Student Book 2 Reference⟩ [参考資料] Evaluation Criteria⟩ Quiz 1 25%; Skit 1 25%; Quiz 2 25%; Skit 1 25% Re-evaluation⟩ [再評価] Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220911 Contact (Office-Hour, Room, E-mail)⟩ ⇒ Luxton .